OpenStudy (anonymous):

Please help: Eliminate the parameter. x = 5t, y = t + 8

OpenStudy (anonymous):

x = 5t, y = t + 8 t=x/5, y=(x/5)+8,

OpenStudy (anonymous):

Thanks kitty! Wow youre fast lol

OpenStudy (anonymous):

u r welcome

OpenStudy (anonymous):

I have 1 more its kind of hard. should i make a new question?

OpenStudy (anonymous):

y